# Massive Memorization with Hundreds of Trillions of Parameters for Sequential Transducer Generative Recommenders

### v1

@@ 2 Related Work
The application of transformer architectures in recommendation systems has been explored extensively. By leveraging the self-attention mechanism, transformers can model complex user-item interactions over time, facilitating more nuanced and personalized recommendations (Subbiah and Aggarwal, 2024). The Deep Interest Network (DIN) (Zhou et al., 2018) and its follow-up work, Search-based Interest Modeling (SIM) (Pi et al., 2020; Chang et al., 2023; Si et al., 2024), leverage lifelong sequential behavior data. This approach employs search-based mechanisms, also known as General Search Units (GSUs), to select a small subset of relevant interactions from the user’s historical sequence that are pertinent to the target item followed by a standard transformer network, referred to as Exact Search Units (ESUs), to compute the final target item representation. Notably, this method enables the modeling of user behavior data with lengths reaching up to hundreds of thousands (Pi et al., 2020).
@@
Apart from Flash Attention (Dao et al., 2022; Dao, 2023) that is designed to improve the efficiency of the softmax attention mechanisms, there is a new trend to explore linear complexity attention mechanisms. Katharopoulos et al. (2020a) first proposes linear attention. By applying matrix multiplication associative property, it enables a change in computation order from $(QK^{T})V$ to $Q(K^{T}V)$, reducing computation complexity from $O(N^{2})$ to $O(N)$ with respect to sequence length $N$. Recently, Lightning Attention v1 (Qin et al., 2024a) and v2 (Qin et al., 2024b) propose a light network which contains Gated Linear Attention (GLA) and Simple Gated Linear Unit (SGLU) to make linear attention more practical. Another branch of linear complexity work, namely state space model (SSM), has been widely studied. Mamba (Gu and Dao, 2024) is a pioneering work in SSM and widely used in many real-world applications, followed by Hydra (Hwang et al., 2024) which is the double-headed version of Mamba to address non-causal scenarios.
@@ 3.2.1 Linear Attention with Candidate Items for Recommendation
With the emergence of Large Language Models (LLMs), researchers have proposed some linear complexity attention algorithms to accelerate transformer blocks (Katharopoulos et al., 2020a; Qin et al., 2024a; b; Han et al., 2024). However in recommendation systems, unlike the text sequences in LLM, a strict rule is that the candidates cannot attend each other, since it introduces label leakage due to the fact that the logged candidates typically only form a small subset of the input candidates during inference. Therefore, we propose a linear-complexity self attention mechanism that avoids attention among candidates.
@@
where $Q$, $K$ and $V$ have shape $(L,d)$ and $L$ is the sequence length. Then the original linear attention (Katharopoulos et al., 2020b) for a UIH sequence $S$ can be written similarly as follows
@@
Note that candidates cannot attend to each other. This is a strict rule in recommendation systems otherwise the model training will fail due to the leakage between candidate items. It gets slightly trickier if we also want each candidate to attend to itself. Instead of $TT^{\top}T$, the contribution due to the self attention of each target item to herself is given by
@@ 3.2.2 Quasi-linear Attention for Recommendation
The quasi-linear attention contains two parts: Quasi Linear Unit (QLU) module and Simple Gated Linear Unit (SGLU) module. The QLU module aims to model the interaction of $Q,K,V$ matrices with SiLU non-linear activation as shown in Figure LABEL:fig:Quasi-la. For the SGLU module, we use the same gated function as TransNormerLLM (Qin et al., 2024a).
@@ 3.2.3 Generative Sequence Reconstruction Loss
Since causal transformer block ensures that the output embedding $v_{i}$ only depends on $u_{1},\ldots,u_{i}$, there is no leak of information from $u_{i+1}$ to $v_{i}$. This forces the personalized seed embeddings $s_{i}$ to maximize information retained of the user history sequence $u_{1},\ldots,u_{M}$. Similar ideas have roots in the Variational Auto-Encoder (Kingma and Welling, 2022), and have appeared in the context of transformer networks recently (Henderson and Fehr, 2022). However to the best of our knowledge, there has not been any explicit use in recommendation.
@@ 4 Embedding Delivery System
Figure 6 outlines the system’s end-to-end architecture, which comprises three main stages: (1) online training of the source model using training data stream, (2) delivery of sequence summarization embeddings to downstream models via two routes: a real-time message queue, e.g., Kafka (Kreps et al., 2011) and persistent storage, e.g., Hive (Thusoo et al., 2009), and (3) serving embeddings through a geographically replicated in-memory key-value store. This design ensures both real-time performance and scalability for industrial applications. For scalability, we deliberately compress the user interaction history sequence to $O(100)$ terabytes level, making it feasible to deploy to existing systems.
@@ 5.1.2 Baselines and Evaluation Metrics
We use normalized entropy (NE) (He et al., 2014) as our evaluation metric, which calculates the cross entropy between the predicted probabilities and the labels, then normalizes it by the entropy of the constant predictor at label average. We also report the area under curve (AUC) for the traditional setting. Additional details about this section are in Appendix C.
Table 2: Comparisons on public and Minimal Production datasets. VISTA-w/-QLA and VISTA-w/o-QLA are the VISTA model with and without quasi-linear attention, respectively.
@@ 5.3 Online A/B Experimental Results
We conducted an online A/B test on our production video recommendation system, using 5% of the entire site traffic during a period of 15 days. The baseline is the HSTU model and we compare with adding the VISTA module, which is the same as our offline experiment setup. Online metrics for the main consumption task “C-Task” and other online onboarding metrics, “O1-Task” and “O2-Task”, were significantly improved by $0.5\%$, $2\%, 0.0$$04\%$, respectively. With a $0.01\%$ “O2-Task” gain considered a substantial improvement on our platform, the VISTA model made realized contributions to the recommendation system.
@@ 6 Conclusion and Discussions
In this paper, we have proposed the VIrtual Sequential Target Attention (VISTA) framework, a novel two-stage approach that compresses ultra-long user interaction histories into a set of compact embeddings. This design strikes a crucial balance between computational efficiency and predictive accuracy, addressing the latency and scalability challenges of processing ultra-long user sequence data in production systems. VISTA’s practical applicability is underscored by its resilience to slight de-synchronization between its stages and its ability to approximate complex transformer architectures without their substantial computational cost. Our empirical evaluations demonstrate that VISTA not only captures the core information within user interactions but also achieves significant improvements across platform metrics. Our plans for future research involve further optimizing VISTA’s compression techniques and exploring its applicationa cross other domains to enhance its generalizability.
7 Acknowledgement
This work results from a large cross organization collaboration. It would not be possible without contributions from the collaborators and supports from the leaderships as follows (alphabetic order): Zheng-Yong Ang, David Bauer, Connor Chen, Siqiao Chen, Huihui Cheng, Ek Kheng Chung, Litao Deng, Litao Deng, Shilin Ding, Chenhao Feng, Kevin Goulding, Liang Guo, Mengyue Hang, Chufeng Hu, Jizhou Huang, Yanzun Huang, Han Jiang, Justin Khim, Emma Lin, Yang Liu, Li Lu, Wenhan Lyu, Jing Ma, Matt Ma, Jing Qian, Yongxiong Ren, Xinyue Shen, Daisy Shi, Hongzheng Shi, Ge Song, Yisong Song, Wanting Tan, Hao Wan, Meihong Wang, Hong Yan, Yihang Yang, Christina You, Rui Zhang, Yue Zhang, John Zheng, Lizhen Zhu.
@@ B.2 Activation for Quasi-Linear Attention
Appendix C Additional Experiment Details
C.1 Datasets
@@
The KuaiRand dataset by Gao et al. (2022) is a sequential recommendation dataset collected from the recommendation logs of the video-sharing mobile app Kuaishou. The KuaiRand-1K subset contains a random sample 1,000 users after removing irrelevant videos. There are 4 million videos remaining in this subset. Our experiments use a subset of all features available in KuaiRand-1K, namely user id, video is, video is history, click history, like history, and lvv (long video view) history.
@@ C.2 FuxiCTR Framework
C.3 More on VISTA’s Two-Stage Attention

### v2

@@ 2 Related Work
The application of transformer architectures in recommendation systems has been explored extensively. By leveraging the self-attention mechanism, transformers can model complex user-item interactions over time, facilitating more nuanced and personalized recommendations (Subbiah and Aggarwal, 2024). The Deep Interest Network (DIN) (Zhou et al., 2018) and its follow-up work, Search-based Interest Modeling (SIM) (Pi et al., 2020; Chang et al., 2023; Si et al., 2024), leverage lifelong sequential behavior data. This approach employs search-based mechanisms, also known as General Search Units (GSUs), to select a small subset of relevant interactions from the user’s historical sequence that are pertinent to the target item followed by a standard transformer network, referred to as Exact Search Units (ESUs), to compute the final target item representation. Notably, this method enables the modeling of user behavior data with lengths reaching up to hundreds of thousands (Pi et al., 2020). Other methods (Liu et al., 2023) preprocess user histories into groups and attend to the group embeddings, and separately attend to subsequences in the user history relevant to the target item.
@@
Apart from Flash Attention (Dao et al., 2022; Dao, 2023) that is designed to improve the efficiency of the softmax attention mechanisms, there is a new trend to explore linear complexity attention mechanisms. Katharopoulos et al. (2020) first proposes linear attention. By applying matrix multiplication associative property, it enables a change in computation order from $(QK^{T})V$ to $Q(K^{T}V)$, reducing computation complexity from $O(N^{2})$ to $O(N)$ with respect to sequence length $N$. Recently, Lightning Attention v1 (Qin et al., 2024a) and v2 (Qin et al., 2024b) propose a light network which contains Gated Linear Attention (GLA) and Simple Gated Linear Unit (SGLU) to make linear attention more practical. Another branch of linear complexity work, namely state space model (SSM), has been widely studied. Mamba (Gu and Dao, 2024) is a pioneering work in SSM and widely used in many real-world applications, followed by Hydra (Hwang et al., 2024) which is the double-headed version of Mamba to address non-causal scenarios.
@@ 3.2.1 Linear Attention with Candidate Items for Recommendation
With the emergence of Large Language Models (LLMs), researchers have proposed some linear complexity attention algorithms to accelerate transformer blocks (Katharopoulos et al., 2020; Qin et al., 2024a; b; Han et al., 2024). However in recommendation systems, unlike the text sequences in LLM, a strict rule is that the candidates cannot attend each other, since it introduces label leakage due to the fact that the logged candidates typically only form a small subset of the input candidates during inference. Therefore, we propose a linear-complexity self attention mechanism that avoids attention among candidates.
@@
where $Q$, $K$ and $V$ have shape $(L,d)$ and $L$ is the sequence length. Then the original linear attention (Katharopoulos et al., 2020) for a UIH sequence $S$ can be written similarly as follows
@@
Note that candidates cannot attend to each other. This is a strict rule in recommendation systems otherwise the model training will fail due to the leakage between candidate items. It gets slightly trickier if we also want each candidate to attend to itself. Instead of $TT^{\top}T$, the contribution due to the self attention of each target item to itself is given by
@@ 3.2.2 Quasi-linear Attention for Recommendation
The quasi-linear attention contains two parts: Quasi Linear Unit (QLU) module and Simple Gated Linear Unit (SGLU) module. The QLU module aims to model the interaction of $Q,K,V$ matrices with SiLU non-linear activation as shown in Figure 5. For the SGLU module, we use the same gated function as TransNormerLLM (Qin et al., 2024a).
@@ 3.2.3 Generative Sequence Reconstruction Loss
Since causal transformer block ensures that the output embedding $v_{i}$ only depends on $u_{1},\ldots,u_{i}$, there is no leak of information from $u_{i+1}$ to $v_{i}$. This forces the personalized seed embeddings $s_{i}$ to maximize information retained of the user history sequence $u_{1},\ldots,u_{M}$. Similar ideas have roots in the Variational Auto-Encoder (Kingma and Welling, 2022), and have appeared in the context of transformer networks recently (Henderson and Fehr, 2022). However to the best of our knowledge, there has not been any explicit use in recommendation. For more discussion on this reconstruction loss, see Appendix C.
@@ 4 Embedding Delivery System
Figure 6 outlines the system’s end-to-end architecture, which comprises three main stages: (1) online training of the source model using training data stream, (2) delivery of sequence summarization embeddings to downstream models via two routes: a real-time message queue, e.g., Kafka (Kreps et al., 2011) and persistent storage, e.g., Hive (Thusoo et al., 2009), and (3) serving embeddings through a geographically replicated in-memory key-value store. In our system, we update the summarization embeddings on a 2-hour cadence, which was shown to have similar performance compared to using the summarization module directly in online A/B tests. This design ensures both real-time performance and scalability for industrial applications. For scalability, we deliberately compress the user interaction history sequence to $O(100)$ terabytes level, making it feasible to deploy to existing systems.
@@ 5.1.2 Baselines and Evaluation Metrics
We use normalized entropy (NE) (He et al., 2014) as our evaluation metric, which calculates the cross entropy between the predicted probabilities and the labels, then normalizes it by the entropy of the constant predictor at label average. We also report the area under curve (AUC) for the traditional setting. Additional details about this section are in Appendix D.
Table 2: Comparisons on public and Minimal Production datasets. VISTA-w/-QLA and VISTA-w/o-QLA are the VISTA model with and without quasi-linear attention, respectively. [^5] Comparisons to the next best model are significant at the $p<0.001$ level when using a paired t-test. ANOVA tests with multiple next-best models are significant at the $p<0.001$ level.
@@ 5.3 Online A/B Experimental Results
We conducted an online A/B test on our production video recommendation system, using 5% of the entire site traffic during a period of 15 days. The baseline is the HSTU model and we compare with adding the VISTA module, which is the same as our offline experiment setup. Online metrics for the main consumption task “C-Task” and other online onboarding metrics, “O1-Task” and “O2-Task”, were significantly improved by $0.5\%$, $2\%, 0.0$$04\%$, respectively. VISTA demonstrated a 94% reduction in inference GPU resource (measured by inference QPS) usage by caching and serving embeddings, rather than re-computing them for every new user request. With a $0.01\%$ “O2-Task” gain considered a substantial improvement on our platform, the VISTA model made realized contributions to the recommendation system.
@@ 6 Conclusion and Discussions
In this paper, we have proposed the VIrtual Sequential Target Attention (VISTA) framework, a novel two-stage approach that compresses ultra-long user interaction histories into a set of compact embeddings. This design strikes a crucial balance between computational efficiency and predictive accuracy, addressing the latency and scalability challenges of processing ultra-long user sequence data in production systems. VISTA’s practical applicability is underscored by its resilience to slight de-synchronization between its stages and its ability to approximate complex transformer architectures without their substantial computational cost. Our empirical evaluations demonstrate that VISTA not only captures the core information within user interactions but also achieves significant improvements across platform metrics. Our plans for future research involve further optimizing VISTA’s compression techniques and exploring its applications across other domains to enhance its generalizability.
Acknowledgement
This work results from a large cross organization collaboration. It would not be possible without contributions from the collaborators and supports from the leaderships as follows (alphabetic order): Zheng-Yong Ang, David Bauer, Connor Chen, Shouwei Chen, Siqiao Chen, Huihui Cheng, Ek Kheng Chung, Litao Deng, Shilin Ding, Chenhao Feng, Kevin Goulding, Liang Guo, Mengyue Hang, Maxwell Lin-He, Xiaoxin He, Chufeng Hu, Jizhou Huang, Yanzun Huang, Han Jiang, Justin Khim, Emma Lin, Zihan Li, Yang Liu, Yining Liu, Li Lu, Wenhan Lyu, Jing Ma, Matt Ma, Jing Qian, Rui Qiao, Chuyu Qiu, Yongxiong Ren, Xinyue Shen, Daisy Shi, Hongzheng Shi, Ge Song, Yisong Song, Wanting Tan, Hao Wan, Meihong Wang, Yanhong Wu, Hong Yan, Yihang Yang, Chuanwei Yi, Christina You, Haoli Zhang, Rui Zhang, Yue Zhang, John Zheng, Xinye Zheng, Lizhen Zhu, Maggie Zhuang.
@@ B.2 Activation for Quasi-Linear Attention
Appendix C More on Reconstruction Loss
The reconstruction loss is a measure of how much information of the full UIH is captured by VISTA’s virtual embeddings. We used L2 norm to measure how much we can reconstruct the original UIH given the virtual embedding as input, and we have verified that it is an informative metric for us to quantitatively measure the reconstruction quality.
Figure 11: Comparing the reconstruction loss over training steps with and without explicitly including it in the total loss on the Amazon dataset. The dotted line is the final evaluation reconstruction loss.
On the Amazon dataset, for example, we can see that training the VISTA model without explicitly minimizing the reconstruction loss still reduces the reconstruction loss of the learned embedding against the full UIH as the model improves. However, the reconstruction loss plateaus and the model takes longer to converge after some training steps (as our training pipeline supports early stopping). With the explicit introduction of the reconstruction loss, we see a dramatic decrease in the reconstruction loss in the first training steps and faster model convergence. The test metrics also improved by 0.22% AUC and 1.11% NE with the use of the reconstruction loss.
Appendix D Additional Experiment Details
D.1 Datasets
@@
The KuaiRand dataset by Gao et al. (2022) is a sequential recommendation dataset collected from the recommendation logs of the video-sharing mobile app Kuaishou. The KuaiRand-1K subset contains a random sample 1,000 users after removing irrelevant videos. There are 4 million videos remaining in this subset. Our experiments use a subset of all features available in KuaiRand-1K, namely user id, video id, video id history, click history, like history, and lvv (long video view) history.
@@ D.2 FuxiCTR Framework
D.3 More on VISTA’s Two-Stage Attention